​If the reserve ratio increased from 5 percent to 10 percent, then the money multiplier would

  a.

​rise from 5 to 10.

 
  b.

​rise from 10 to 20.

 
  c.

​fall from 20 to 10.

 
  d.

​fall from 10 to 5.
Answer #1

Answer

option c

Multiplier =1/reserve ratio

rr=5% then multiplier =1/0.05=20

rr=10% then multiplier =1/0.1=10

the multiplier falls from 20 to 10 as the reserve ratio increases because there is an inverse relationship between rr and multiplier.